# thanks: Full author list given at the end of the paper

### v2

@@ IV.2 Decay asymmetry and SM test
$\langle\Delta_{BP}\rangle$ $0.637\pm 0.444\pm 0.014$ $2.190\pm 0.730\pm 0.029$ $1.901\pm 0.603\pm 0.040$
$\langle\beta_{BP}\rangle$ $0.365^{+0.173}_{-0.246}\pm 0.010$ $0.704^{+0.143}_{-0.480}\pm 0.015$ $0.764^{+0.051}_{-0.237}\pm 0.018$
$\langle\gamma_{BP}\rangle$ $0.637^{+0.103}_{-0.202}\pm 0.011$ $0.502^{+0.591}_{-0.303}\pm 0.021$ $0.262^{+0.478}_{-0.383}\pm 0.031$
$\delta_{p}-\delta_{s}$ $2.71^{+0.28}_{-0.17}\pm 0.02$ $19^{+0.49}_{-0.13}\pm 0.02 2.2$$23^{+0.19}_{-0.06}\pm 0.03$
$A^{\alpha_{BP}}_{C\!P}$ $079^{+0.115}_{-0.101}\pm 0.019 0.0$$002\pm 0.047\pm 0.017$ $206^{+0.188}_{-0.156}\pm 0.028 0.0$$086\pm 0.081\pm 0.085$
$\text{tan}\phi_{C\!P}$ $0.232\pm 0.242\pm 0.025$ $393\pm 0.651\pm 0.042 0.0$$007\pm 0.474\pm 0.034$
@@ V Discussion
The effect of transverse polarization of $\Lambda_{c}^{+}$ baryons is discovered for the first time by analyzing $6.4~\text{fb}^{-1}$ of $e^{+}e^{-}$ annihilation data collected at the CM energies between $4600$ MeV and $4951$ MeV with the BESIII detector. The non-zero transverse polarization indicates the phase shift of the non-zero EMFF. For the first time, we scanned the phase shifts at multiple energy points and demonstrated their differences from theoretical predictions. In 2021, Ref. [76] used modified vector meson dominance (VMD) model considering vector charmed mesons and their excitations $J/\psi$, $\psi(2S)$, $\psi(3770)$, $\psi(4040)$, $\psi(4160)$, and $\psi(4415)$. They then constructed and fitted the Born cross section of $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$, since it was studied in 2021, where the data they fitted comes from the Belle Collaboration [79] and the BESIII Collaboration [66]. As the Belle Collaboration provided the main lineshape, the fitting is dominated by Belle’s results. The fitting results well describe the Born cross section of $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$. However, Belle’s results showed a significant enhancement of $Y(4630)$, which was refuted by BESIII’s more precise results fifteen years later [78]. Based on this new result, another work [77] also based on the VMD model using resonances, $\psi(4500)$, $\psi(4660)$, $\psi(4790)$, and $\psi(4900)$, describes the Born cross section very well. Due to higher precision and the absence of abnormal $Y(4630)$ contribution, the fitted results describe the oscillation behavior of $|G_{E}/G_{M}|$ with energy points measured by BESIII Collaboration. Unfortunately, the results of these two works [76, 77] cannot describe the phase shifts between $G_{E}$ and $G_{M}$ very well. The theoretical predictions still show a significant deviation from the experimental measurements. Therefore, the stucture of $\Lambda_{c}^{+}$ and the mechanism of its production in the $e^{+}e^{-}$ collider remains a mystery.
@@
The transverse polarization of the $\Lambda_{c}^{+}$ baryon provides an opportunity to measure the decay asymmetry parameters of the $\Lambda_{c}^{+}$ decay. The golden two-body decays $\Lambda_{c}^{+}\to pK^{0}_{S}$, $\Lambda\pi^{+}$, $\Sigma^{0}\pi^{+}$, and $\Sigma^{+}\pi^{0}$ have been studied. The parameter $\alpha_{pK^{0}_{S}}$ has changed its sign, replacing the previous results [15]. The value of $\alpha_{pK^{0}_{S}}$ is now consistent with most theoretical calculations. The decay asymmetry parameters $\alpha_{\Sigma^{+}\pi^{0}}$ and $\alpha_{\Sigma^{0}\pi^{+}}$ support the prediction under $SU(3)$ flavor symmetry but $\alpha_{\Sigma^{+}\pi^{0}}$ is more than $2\sigma$ different than the measurements from Belle [27]. Due to the cascading decay in $\Lambda_{c}^{+}\to\Lambda\pi^{+},~\Sigma^{0}\pi^{+}$, and $\Sigma^{+}\pi^{0}$, we also obtain the value of $\Delta_{BP}$, and use this to derive $\beta_{BP}$ and $\gamma_{BP}$. This also enables us to extract the $S$- and $P$-wave phase shift, $\delta_{p}-\delta_{s}$. The results of phase shifts in this work are consistent with recent predictions [30, 31, 33] following the derived method from the LHCb Collaboration [32], and we use the convention that $sign=+1$ as in Ref. [14]. We have also derived the dynamic parameters $|A|$ and $|B|$ for weak interactions. Previous theoretical studies on the decay of charmed baryons have also predicted $|A|$ and $|B|$, but most of them did not account for the contribution of phase shifts to the decay asymmetry parameters, which complicates the accurate prediction of these parameters. Only a few recent works have emphasized the impact of phase shifts and predicted $|A|$ and $|B|$ [30, 31, 33]. Clearly, the reliability of their predictions is improving. The results presented here supersede those in Refs [15, 78].
@@
A strong phase shift can be significant in physical processes and theoretically can greatly improve the CPC\!P asymmetry prediction [80, 29, 31]. In our work, we calculated the $\delta_{p}-\delta_{s}$ in three channels and demonstrated the existence of a large phase shift in charmed baryon decays. Such a large phase shift can be induced by re-scattering processes and loop effects, and further consideration of the role of final state interactions is warranted [29]. The decays measured in this work are all CF processes, where weak phase shifts are zero. Although the CPC\!P asymmetry phenomenon is impossible in the CF processes, it can still be transferred to other Cabibbo-suppressed processes through symmetry in physics. As discussed in Ref. [80], the re-scattering effects can enhance CPC\!P asymmetry by an order of magnitude, which makes it promising to observe CPC\!P asymmetry effect in baryon decays. In addition, based on the separate measurements of the positive and negative processes, we also test $A^{\alpha_{BP}}_{C\!P}$, strong and weak phase shifts. They are related by $A^{\alpha_{BP}}_{C\!P}=-\tan(\delta_{P}-\delta_{S})\tan(\xi_{P}-\xi_{S})$, and the results show that $A^{\alpha_{BP}}_{C\!P}$ and weak phase shifts are consistent with zero, which aligns with the expectations of the SM.
@@
In conclusion, the observation of the transverse polarization effect of the $\Lambda_{c}^{+}$ with the BESIII detector represents a complete description in charmed baryon production in $e^{+}e^{-}$ annihilations, contributing to our understanding of its production properties. Our analysis of decay asymmetry parameters for $\Lambda_{c}^{+}$ decays into $pK^{0}_{S}$, $\Lambda\pi^{+}$, $\Sigma^{0}\pi^{+}$, and $\Sigma^{+}\pi^{0}$ not only updates and validates previous measurements but also reveals significant strong phase shifts, providing valuable constraints on various theoretical models. Furthermore, the systematic study of production and decay properties provides a comprehensive dataset that refines global fits in flavor physics, enhances the understanding of CPC\!P violation in the charm sector, and aligns with current SM predictions. These findings, combined with the precise measurement of polarization dynamics, open new avenues for investigating hadronic interactions, spin effects, and charm quark dynamics. This work establishes a robust framework for future experimental programs at Belle II, LHCb, STCF and other facilities, setting the stage for deeper insights into the non-perturbative aspects of QCD and the exploration of potential new physics.

### v3

@@ IV.2 Decay asymmetry and SM test
$\langle\Delta_{BP}\rangle$ $\phantom{-}0.637\pm 0.444\pm 0.014$ $\phantom{-}2.190\pm 0.730\pm 0.029$ $\phantom{-}1.901\pm 0.603\pm 0.040$
$\langle\beta_{BP}\rangle$ $\phantom{-}0.365^{+0.173}_{-0.246}\pm 0.010$ $\phantom{-}0.704^{+0.143}_{-0.480}\pm 0.015$ $\phantom{-}0.764^{+0.051}_{-0.237}\pm 0.018$
$\langle\gamma_{BP}\rangle$ $\phantom{-}0.637^{+0.103}_{-0.202}\pm 0.011$ $0.502^{+0.591}_{-0.303}\pm 0.021$ $0.262^{+0.478}_{-0.383}\pm 0.031$
$\delta_{p}-\delta_{s}$ $\phantom{-}2.71^{+0.28}_{-0.17}\pm 0.02$ $\phantom{-}2.19^{+0.49}_{-0.13}\pm 0.02$ $\phantom{-}2.23^{+0.19}_{-0.06}\pm 0.03$
$A^{\alpha_{BP}}_{C\!P}$ $\phantom{-}0.079^{+0.115}_{-0.101}\pm 0.019$ $\phantom{-}0.002\pm 0.047\pm 0.017$ $\phantom{-}0.206^{+0.188}_{-0.156}\pm 0.028$ $0.086\pm 0.081\pm 0.029$
$\text{tan}\phi_{C\!P}$ $\phantom{-}0.232\pm 0.242\pm 0.025$ $\phantom{-}0.393\pm 0.651\pm 0.042$ $0.007\pm 0.474\pm 0.034$
@@ V Discussion
The effect of transverse polarization of $\Lambda_{c}^{+}$ baryons is discovered for the first time by analyzing $6.4~\text{fb}^{-1}$ of $e^{+}e^{-}$ annihilation data collected at the CM energies between $4600$ MeV and $4951$ MeV with the BESIII detector. The observed transverse polarization reveals that the electric and magnetic form factors of the $\Lambda^{+}_{c}$ carry a relative phase, reflecting complex dynamics in its electromagnetic structure. For the first time, we further mapped how this phase varies with collision energy, revealing clear deviations from theoretical expectations.
The vector meson dominance (VMD) model offers a framework to describe how the production cross sections of the baryon-antibaryon pairs via electron-positron annihilations change with energy [Li:2021lvs, Chen:2023oqs, Chen:2024luh], assuming that the initial electromagnatic process dominates with virtual photon coupling to the produced hadrons primarily through intermediate vector mesons. This approach has also been successfully applied to describe nucleon form factors [Yang:2019mzq]. In 2021, a VMD model incorporating the vector charmonium states was proposed [Wan:2021ncg] to describe the Born cross sections of $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ measured from Belle [Belle:2008xmh] and BESIII [BESIII:2017kqg]. That time Belle data determined the overall lineshape, and the form factors $|G_{E}/G_{M}|$ in the VMD model were less constrained with sparse BESIII data. Two years later, BESIII reported the cross sections and the form factor ratio $|G_{E}/G_{M}|$ with much improved precisions at more energy points [BESIII:2023rwv], which constrains the lineshape largely. A subsequent VMD-based study including higher vector states such as $\psi(4500)$, $\psi(4660)$, $\psi(4790)$, and $\psi(4900)$ is drived to achieve an excellent description of the cross sections and $|G_{E}/G_{M}|$ at different energy points [Chen:2023oqs]. Unfortunately, neither calculation [Wan:2021ncg, Chen:2023oqs] reproduces the measured phase difference between $G_{E}$ and $G_{M}$, indicating a persistent discrepancy between theory and experiment. This suggests that the internal structure of the $\Lambda_{c}^{+}$ and its production dynamics in $e^{+}e^{-}$ annihilation are not yet fully understood.
@@
The observed transverse polarization enables precise measurements of the $\Lambda_{c}^{+}$ decay asymmetry parameters [Li:2025nzx]. The key two-body decays $\Lambda_{c}^{+}\to pK^{0}_{S}$, $\Lambda\pi^{+}$, $\Sigma^{0}\pi^{+}$, and $\Sigma^{+}\pi^{0}$ have been studied. The measured $\alpha_{pK^{0}_{S}}$ now changes the sign of center value and is consistent with most theoretical calculations. The decay asymmetry parameters $\alpha_{\Sigma^{+}\pi^{0}}$ and $\alpha_{\Sigma^{0}\pi^{+}}$ agree with the calculation under $SU(3)$ flavor symmetry but $\alpha_{\Sigma^{+}\pi^{0}}$ is more than $2\sigma$ different than the measurements from Belle [Belle:2022bsi]. Due to the cascading decay in $\Lambda_{c}^{+}\to\Lambda\pi^{+},~\Sigma^{0}\pi^{+}$, and $\Sigma^{+}\pi^{0}$, we also obtain the value of $\Delta_{BP}$, and use this to derive $\beta_{BP}$. This enables us to extract the $S$- and $P$-wave phase shift, $\delta_{p}-\delta_{s}$. The results of phase shifts in this work are consistent with recent theoretical calculations [Zhong:2024zme, Zhong:2024qqs, Cheng:2024lsn] under the convention $sign=+$ as proposed in Ref. [Wang:2024wrm]. The extra parameter signsign is introduced in the unified parameterization of $S$- and $P$-wave amplitudes in baryon decays and is important when conducting global analyses of baryon polarization, involving the strong phase shifts, among different experiments and performing future searches for baryon CPC\!P violation [Wang:2024wrm]. We have also derived the dynamic parameters $|A|$ and $|B|$ for weak interactions. Previous theoretical studies on the decay of charmed baryons have also calculated $|A|$ and $|B|$, but most of them did not account for the contribution of phase shifts to the decay asymmetry parameters, which complicates the accurate calculation of these parameters. Only a few recent works have emphasized the impact of phase shifts and predicted $|A|$ and $|B|$ [Zhong:2024zme, Zhong:2024qqs, Cheng:2024lsn]. Clearly, the reliability of their calculations is improving. The results presented here supersede those in Refs. [BESIII:2019odb, BESIII:2023rwv].
@@
A strong phase shift can be significant in physical processes and theoretically can greatly improve the CPC\!P asymmetry prediction [He:2024unv, Geng:2023pkr, Zhong:2024qqs]. In our work, we calculate $\delta_{p}-\delta_{s}$ in three channels and demonstrate the existence of a large phase shift in charmed baryon decays, as reflected with the values of $\tan\Delta_{S}$ in Table 4. Such a large phase shift can be induced by re-scattering processes and loop effects [Geng:2023pkr, Jia:2024pyb], hence indicating that final-state interactions play an important role in QCD calculations. The decays measured in this work are all CF processes, where weak phase shifts are zero. Although the CPC\!P asymmetry phenomenon is impossible in the CF processes, it can still be transferred to other Cabibbo-suppressed processes through symmetry in physics. As discussed in Ref. [He:2024unv], the re-scattering effects can enhance CPC\!P asymmetry by an order of magnitude, which makes it promising to observe CPC\!P asymmetry effect in baryon decays. In addition, based on the separate measurements of the positive and negative processes, we also test $A^{\alpha_{BP}}_{C\!P}$, strong and weak phase shifts. They are related by $A^{\alpha_{BP}}_{C\!P}=-\tan(\delta_{p}-\delta_{s})\tan(\xi_{p}-\xi_{s})$, and the results show that $A^{\alpha_{BP}}_{C\!P}$ and weak phase shifts are consistent with zero, which aligns with the expectations of the SM.
@@
In conclusion, the observation of the transverse polarization effect of the $\Lambda_{c}^{+}$ with the BESIII detector represents a complete description in charmed baryon production in $e^{+}e^{-}$ annihilations, contributing to our understanding of its production properties. Our analysis of decay asymmetry parameters for $\Lambda_{c}^{+}$ decays into $pK^{0}_{S}$, $\Lambda\pi^{+}$, $\Sigma^{0}\pi^{+}$, and $\Sigma^{+}\pi^{0}$ not only updates and validates previous measurements but also reveals significant strong phase shifts, providing valuable constraints on various theoretical models. Furthermore, the systematic study of production and decay properties provides a comprehensive dataset that refines global fits in flavor physics, enhances the understanding of QCD in charm sector and the capability of prediction of the CPC\!P asymmetry in charmed baryons. These findings, combined with the precise measurement of polarization dynamics, open new avenues for investigating hadronic interactions, spin effects, and charm quark dynamics. This work establishes a robust framework for future experimental programs at Belle II, LHCb, STCF [Wang:2025yjs] and other facilities, setting the stage for deeper insights into the non-perturbative aspects of QCD and better sensitivities in searching for CPC\!P violations, as well as the exploration of potential new physics.